OpenStudy (anonymous):

5(x-1) over 4 equals 3(x+1) over 2. is x=2?

OpenStudy (anonymous):

1/4 does *not* equal 9/2

OpenStudy (anonymous):

you mean 5/4 doesnt equal 9/2

OpenStudy (anonymous):

\[\frac{5(x-1)}{4}=\frac{3(x+1)}{2}\]so,\[5(x-1)=6(x+1)\]and,\[5x-5=6x+1\]giving,\[x=-6\]

OpenStudy (anonymous):

ooops 5x-5=6x+6 x=-11

OpenStudy (mertsj):